OpenStudy (anonymous):

Solve –2(x – 1) = 18

OpenStudy (anonymous):

1-2x=18 1-18=2x x=-17/2

jimthompson5910 (jim_thompson5910):

-2(x-1)=18 -2x+2=18 -2x=18-2 -2x=16 x=16/(-2) x=-8